If cos x=4/5 and x is an acute angle, find the exact value of cos2x, in simplest fractional form
Answer by mananth(16946)

cos(2x) = 2 cos^2x - 1
cos x = 4/5
cos^2x = 16/25
2*cos^2x= 32/25
2*cos^2x-1= 32/25 -1
=7/25
cos(2x)= 7/25